It has been alot of years, can you tell me?

1 = 1+ d\33 how do I solve for d.

2 = 1+ d\33 etc.
do not remember how do get rid of the one on the right side of the equation and to get rid of the /33 if I have to multiply it by the one on the right side as well as the d and the left side as well? tks//mark..

Answer by wuwei96815(245)

1 = 1 + d/33
Multiply all terms by 33
33 = 33 + d
Add -33 to both sides
33 - 33 = 33 - 33 + d
0 = d

2 = 1 + d/33
Multiply all terms by 33
66 = 33 + d
Add -33 to both sides
66 - 33 = 33 - 33 + d
33 = d